# The Structure of the Majorana Clifford Group

### v2

@@ Abstract
In quantum information science, Clifford operators and stabilizer codes play a central role for systems of qubits (or qudits). In this paper, we study the analogous objects for systems of Majorana fermions. A crucial role is played by fermion parity symmetry, which is an unbreakable symmetry present in any system in which the fundamental degrees of freedom are fermionic. We prove that the subgroup of parity-preserving fermionic Cliffords can be represented by the orthogonal group over the binary field $\mathbb{F}_{2}$, and we show how it can be generated by braiding operators and used to construct any (even-parity) Majorana stabilizer code. We also analyze the frame potential for this so-called p-Clifford group, proving that it is equivalent to the frame potential of the ordinary Clifford group acting on a fixed-parity sector of the Hilbert space.
@@ I Introduction
The standard model of quantum computation is formulated in terms of commuting qubit degrees of freedom, but one can also formulate a model of quantum computation in terms of anti-commuting fermionic degrees of freedom [1]. In the standard model, an important role is played by the Clifford group, which is a special subgroup of the full unitary group whose elements act in a simple way on Pauli operators [2]. Here we study the fermionic analog of this structure, the Majorana Clifford group, which is a special subgroup of unitaries whose elements act in a simple way on Majorana fermion operators. And while any Majorana Clifford turns out be just a Jordan-Wigner transformation away from becoming a Pauli Clifford, the extra ingredient of fermion superselection places new constraints on which operations can be considered to be physical.
@@
One motivation for an analysis of the fermionic Clifford group is provided by a promising class of fault-tolerant quantum computer architectures that employs the topological braiding of Majorana fermion modes to encode logical information [3, 4, 5, 6, 7, 8]. While their topological structure provides these systems with some inherent degree of protection against outside noise, the concatenation of additional error-correcting (stabilizer) codes generated by fermionic Clifford operations acting on these Majorana modes will likely still be necessary to achieve adequate macroscopic resilience. Fermionic stabilizer codes, which were initially proposed by Bravyi et al. [9], have since garnered a rich literature [10, 11, 12, 13, 14, 15]. Such codes have advantages compared to their Pauli-based counterparts, most notably their natural protection against parity-switching errors thanks to fermionic superselection rules.
@@
Another motivation arises from the Sachdev-Ye-Kitaev (SYK) model which has recently become one of the most-studied physical systems in the fields of condensed matter and quantum gravity [16, 17, 18, 19, 20]. Describing a large number of randomly coupled Majorana fermions, it manages to provide a simple toy model for holography, connecting concepts in quantum information and high energy physics. Recently, we proposed a tensor network model of the SYK model and analyzed a qubit-based toy model employing random Clifford gates [21]. The desire to generalize this toy model to fermionic degrees of freedom was the starting point of this work, although we have deferred any analysis of the corresponding fermionic tensor networks to future work. Random sampling from the fermionic Clifford group could also be of use for questions concerning fermionic random circuits and fermionic classical shadows [22].
@@
Given these motivations and the broadly important role played by the ordinary Clifford group in quantum information and physics, this paper presents a thorough analysis of the fermionic Clifford group. We are primarily concerned with the description and manipulation of systems composed of $2n$ Majorana fermion modes $\chi_{i}$, $i=1,\ldots,2n$ obeying
@@
and equipped with a fundamental representation on a $2^{n}$-dimensional Hilbert space. Fermion superselection is encoded via their parity symmetry, which is generated by the Hermitian operator
@@
Fermion parity is important because any physical fermionic observable or circuit must commute with it. Hence, we focus on p-Cliffords, the subgroup of Majorana Cliffords which preserve fermion parity. Our findings show that the group of p-Cliffords admits a projective representation as binary orthogonal matrices, a subgroup of the symplectic group describing all Cliffords. Each orthogonal matrix can be generated by a small set of Householder reflections, which in the Hilbert space representation corresponds to the p-Cliffords being generated by Majorana braiding operators. Using a canonical basis for even-parity stabilizers, we show that by, adding ancilla modes, any other even stabilizer basis (and its corresponding logical operators) can be generated by applying an appropriate p-Clifford. Last but not least, we prove that the p-Clifford group can be considered to be a unitary design on a fixed-parity sector of the Hilbert space, but like the ordinary Clifford group only manages to be a 3-design.
Overall, the rest of this paper is structured as follows: In Section II we formally define Majorana strings, prove some basic properties, and discuss the the connections between operator weight, fermion parity and the resulting superselection rules. We also highlight the relation to Pauli strings and the resulting Hilbert (sub-)space representation. Next, in Section III we formally define the general (fermionic) Clifford group, and identify the p-Cliffords as its subgroup, characterized in terms of orthogonal matrices over the field $\mathbb{F}_{2}$ and efficiently generated by braiding operators. In Section IV we discuss fermionic stabilizer states and how, with the help of an ancilla systems, they can be generated using p-Cliffords. Fourth, in Section V we discuss the problem of sampling from the p-Clifford group, analyze the frame potential, and show that the p-Clifford is a 3-design within sectors of fixed fermion parity. We conclude with an outlook and discussion of open problems in Section VI.
@@ II Majorana Strings
To describe strings of fundamental Majorana fermion operators and the superoperators that map between them, it is most convenient to use the language of symplectic projective representations over the binary field $\mathbb{F}_{2}$. This construction follows analogous (and equivalent) programs applied to ordinary Pauli strings [23].
@@ II.1 General Properties
Given $2n$ independent Majorana fermion operators $\chi_{i}$ satisfying (1), we define a general Hermitian Majorana string $\operatorname{\mu}(v)$ as a normal-ordered product of said operators, such that
$\operatorname{\mu}(v)\coloneqq(i)^{v^{T}\operatorname{\operatorname{\omega}_{L}}v}\cdot\chi_{1}^{v_{1}}\chi_{2}^{v_{2}}\cdots\chi_{2n}^{v_{2n}},\,v\in\mathbb{F}_{2}^{2n},$ (3)
@@
It is straightforward to show that this definition indeed satisfies $\operatorname{\mu}(v)^{\dagger}=\operatorname{\mu}(v)$ by reordering the operators after taking the adjoint. All expressions involving elements of the finite field $\mathbb{F}_{2}\equiv(\{0,1\},+,\cdot)$ or vector spaces over $\mathbb{F}_{2}$ are understood to be evaluated mod 2.
With this we can define the equivalent of the Pauli group as the span of all (not necessarily Hermitian) Majorana strings:
@@
The group structure of $\operatorname{\mathcal{M}}_{2n}$ follows directly from the composition property of two (Hermitian) Majorana strings
@@
Here $\zeta(v,v^{\prime})$ is some non-trivial [^1] The reason for these cumbersome expressions is the incompatibility between $\mathbb{F}_{2}$ having addition mod 2 and $(i)^{a}=(i)^{a+4}$ implying addition mod 4. The same issue also plagues qubit Pauli strings. coefficient that can be shown to be of the form
@@
where we also identified the symmetric function $f(v,v^{\prime})=f(v^{\prime},v)$:
@@
Both expressions $\zeta(v,v^{\prime})$ and $f(v,v^{\prime})$ involve a new symmetric [^2] Analogous matrices over fields of higher order are actually skew-symmetric, but for $\mathbb{F}_{2}$ this makes no difference. matrix $\operatorname{\omega}$ which arises in the derivation as the sum of $\operatorname{\operatorname{\omega}_{L}}$ and its transpose $\operatorname{\operatorname{\omega}_{U}}$:
@@
While (7) and (8) are highly nontrivial expressions, it is straightforward to argue (and explicitly prove) that
@@
since $\mu(v)$ is Hermitian with ±$\pm 1$ eigenvalues, and therefore satisfies $\mu(v)^{2}=I\equiv\mu(0)$.
@@
$\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})=(-1)^{v^{T}\operatorname{\omega}v^{\prime}}\operatorname{\mu}(v^{\prime})\operatorname{\mu}(v)$ (13)
The expression $v^{T}\operatorname{\omega}v^{\prime}$ therefore indicates if two strings commute or anti-commute. And since any operator commutes with itself, it also has to satisfy
@@
defines a symplectic product, i.e. it is bilinear, alternating ($\left\langle v,v\right\rangle=0$) and non-degenerate. $\operatorname{\omega}$ is therefore called a symplectic form.
@@
II.3 Weight & Parity
As explained in Section I, two important properties of a Majorana string $\operatorname{\mu}(v)$ (and the associated vector $v$) are its (operator) weight and (fermion) parity.
Definition 4 (Operator Weight & Fermion Parity).
@@ II.3 Weight & Parity
  • The (fermion) parity $\operatorname{p}(\operatorname{\mu}(v))\equiv\operatorname{p}(v)$ of a Majorana string $\operatorname{\mu}(v)$ can be defined as the string’s operator weight mod 2. Equivalently, the parity can be defined in terms of the dot product of the string’s vector representation with itself or $j$ (11) (over $\mathbb{F}_{2}$):
@@
The concept of operator weights is of general interest in the context of stabilizer codes used for quantum error correction. This is because the weight indicates the minimum number of elementary operations one would have to perform in an experimental setup to apply a single Majorana string to a given system. Applying strings with large weights thus requires precise control of the system and a significant amount of time. This allows for errors to occur during the probing, which can therefore be hard to detect. In general, one therefore desires to minimize the weights of the strings being applied as much as possible.
The parity of Majorana strings is of physical relevance due to the constraints imposed on it by fermion superselection. In general, superselection rules dictate that for an observable $O$ (including the Hamiltonian) to be physical, it must obey
$\langle\psi_{1}|O|\psi_{2}\rangle=0$ (18)
for any two states $\lvert\psi_{1}\rangle$, $\lvert\psi_{2}\rangle$ from different superselection sectors. This implies that states from different sectors can’t be put into coherent superpositions, and that the superselection-defining observable is conserved.
In the context of (Majorana) fermions, the two relevant sectors correspond to even and odd parity respectively, and the conserved observable is hence the parity operator
@@
which is uniquely defined up to an overall sign. Superselection then dictates that any (Hermitian) Majorana string describing a physical observable must commute with $(-1)^{F}$. However, using the general property
@@
one comes to the conclusion that $v^{T}\operatorname{\omega}j=v^{T}j=\operatorname{p}(v)$ for all $v\in\mathbb{F}_{2}^{2n}$ per definition (17), and thus the the commutation property (13) implies
@@
Therefore only the strings with even parity commute and hence can act as physical observables or elements of a stabilizer. This does however not imply that odd-parity strings are without use, as they can for example describe logical operators acting on stabilizer states. This is discussed in more detail in Section IV.3.
@@ II.4 Connection to Pauli Strings
$\operatorname{\pi}(v)=(i)^{v^{T}\operatorname{\operatorname{\eta}_{L}}v}\,Z^{v_{1}}X^{v_{2}}\otimes\cdots\otimes Z^{v_{2n-1}}X^{v_{2n}},$ (24)
where $Z$ and $X$ are the standard Pauli matrices and $\operatorname{\operatorname{\eta}_{L}}$ is the equivalent of $\operatorname{\operatorname{\omega}_{L}}$, here corresponding to the lower-triangular part of the matrix
@@
Analogous to $\operatorname{\omega}$, $\eta$ encodes the commutation structure of Pauli strings and hence defines a symplectic product. We can therefore find an invertible linear transformation $W\in\operatorname{GL}(2n,\mathbb{F}_{2})$ that maps between these two symplectic forms:
@@
What $W$ therefore describes is the $\mathbb{F}_{2}$-representation of a Jordan-Wigner transformation between operator product strings of Majorana fermions, and tensor product strings of Pauli operators:
@@
This is trivially true for $n=1$ since in that case $\operatorname{\omega}=\operatorname{\eta}$, which coincides with the natural identification $Z\cong\chi_{1}$, $X\cong\chi_{2}$. In general, $W$ can be shown to be [^3] Plugging (30) into the left side of (26) gives $(\operatorname{\eta}^{c})_{L}\operatorname{\eta}\,(\operatorname{\eta}^{c})_{U}=\operatorname{\eta}\,(\operatorname{\eta}^{c})_{L}(\operatorname{\eta}^{c})_{U}=\operatorname{\eta}\operatorname{\eta}^{c}=\operatorname{\eta}\operatorname{\eta}\operatorname{\omega}=\operatorname{\omega},$ (29) where we used that $(\operatorname{\eta}^{c})_{L}$ commutes with $\operatorname{\eta}$, $\operatorname{\eta}^{c}=(\operatorname{\eta}^{c})_{L}(\operatorname{\eta}^{c})_{U}$, $\operatorname{\eta}^{c}=\operatorname{\eta}\operatorname{\omega}$ because $\operatorname{\eta}$ has odd-parity rows and columns, and $\operatorname{\eta}^{2}=I$.
@@
which is the upper-triangular part of the matrix complement of $\operatorname{\eta}$ (including the diagonal). It is also easy to see that $W$ is its own inverse:
@@
meaning that it maps both vector representations of Majorana strings to ones of Pauli strings, and vice versa. From the structure of (30) it also becomes apparent that low-weight strings in one basis are generally mapped to high-weight strings in the other. This is to be expected, as the most basic Jordan-Wigner transformations are used to assign local fermionic degrees of freedom to topological domain walls in 1D Ising models.

### v3

@@ Abstract
In quantum information science, Clifford operators and stabilizer codes play a central role for systems of qubits (or qudits). In this paper, we study their analogues for systems composed of Majorana fermions. In this case, a crucial role is played by fermion parity symmetry, which is an unbreakable symmetry present in any system with fundamentally fermionic degrees of freedom. We prove that the subgroup of parity-preserving Majorana Cliffords can be represented by the orthogonal group over the binary field $\mathbb{F}_{2}$, and we show how it can be generated by braiding operators and used to construct any (even-parity) Majorana stabilizer code. We also analyze the frame potential for this so-called p-Clifford group when acting on a fixed-parity sector of the Hilbert space, proving that it is equivalent to the frame potential of the ordinary Clifford group acting on the same sector.
@@ I Introduction
The standard model of quantum computation is usually expressed in terms of bosonic (i.e. commuting) degrees of freedom, but one can also formulate an essentially equivalent model in terms of fermionic (i.e. anti-commuting) degrees [1]. In the bosonic formulation, an important role is played by the Clifford group, which is a special subgroup of unitaries whose elements map between tensor products of Pauli operators [2]. Here we study the fermionic analog of this structure, the Majorana Clifford group, whose (unitary) elements act in an equally simple way on operator products of Majorana fermion modes. And while any Majorana Clifford turns out be just a Jordan-Wigner transformation away from becoming a Pauli Clifford, the non-locality of said transformation together with fermion parity superselection places new constraints on which subset of operations can be considered to be physical.
@@
One motivation for an analysis of this fermionic Clifford group is provided by a promising class of fault-tolerant quantum computer architectures that employs the topological braiding of Majorana fermion modes to encode logical information [3, 4, 5, 6, 7, 8]. While their topological structure provides these systems with some inherent degree of protection against outside noise, the concatenation of additional error-correcting (stabilizer) codes generated by Clifford operations acting on these Majorana modes will likely still be necessary to achieve adequate macroscopic resilience. Majorana stabilizer codes, which were initially proposed by Bravyi et al. [9], have since garnered a rich literature [10, 11, 12, 13, 14, 15]. Such codes have advantages compared to their Pauli- and Dirac-based counterparts, most notably their natural protection against errors affecting multiple Majoranas at once, assuming they are sufficiently spatially separated. This is thanks to the aforementioned fermionic superselection rules and the resulting non-locality of observables and parity-preserving operations [16]. The major source of errors for Majorana codes is therefore characterized by single-mode errors, which are referred to as quasiparticle poisoning [12, 15]. Constructing good codes that can detect (and correct) such errors is the subject of ongoing research, to which a systematic analysis of all embedding Majorana Cliffords could prove to be a valuable contribution.
@@
Another motivation arises from the Sachdev-Ye-Kitaev (SYK) model, which has recently become one of the most-studied physical systems in the fields of condensed matter and quantum gravity [17, 18, 19, 20, 21]. Describing a large number of randomly coupled Majorana fermions, it manages to provide a simple toy model for holography, connecting concepts in quantum information and high energy physics. Recently we proposed a potential tensor network ansatz for the SYK model, and analyzed a qudit-based toy example of it by employing random Clifford gates [22]. The desire to generalize this toy model to fermionic degrees of freedom was the starting point of this paper, although we have deferred any analysis of the corresponding fermionic tensor networks to future work.
@@
Given these motivations and the broadly important role played by the ordinary Clifford group in quantum information and physics, this paper presents a thorough analysis of its fermionic counterpart. We are hence primarily concerned with the description and manipulation of systems composed of $2n$ Majorana fermion modes $\chi_{i}$ $(i=1,\ldots,2n)$, obeying
@@
and equipped with a fundamental representation on a $2^{n}$-dimensional Hilbert space. Such systems exhibit a parity symmetry, which is generated by the Hermitian operator
@@
This symmetry imposes constraints on valid physical quantities and operations, which are usually referred to as parity superselection. In general, such superselection rules dictate that for an observable $O$ (including the Hamiltonian) to be physical, it must obey
$\langle\psi_{1}|O|\psi_{2}\rangle=0,$ (3)
for any two states $\lvert\psi_{1}\rangle$, $\lvert\psi_{2}\rangle$ from different superselection sectors (here corresponding to sectors with even and odd parity). This implies that states from different sectors can’t be put into coherent superpositions, and that the superselection-defining observable (in this case $(-1)^{F}$) must be conserved. Hence we primarily focus on p-Cliffords, the subgroup of Majorana Cliffords whose elements commute with $(-1)^{F}$.
Our findings show that the group of p-Cliffords admits a projective representation in terms of binary orthogonal matrices, a subgroup of the symplectic group describing all Cliffords. Each orthogonal matrix can be generated by a small set of Householder reflections, which in the Hilbert space representation corresponds to the p-Cliffords being generated by so-called Majorana braiding operators. Using a canonical basis for even-parity stabilizers, we show that by, adding ancilla modes, any other even stabilizer basis (and its corresponding logical operators) can be generated by applying an appropriate p-Clifford. Last but not least, we prove that the p-Clifford group can be considered to be a unitary design on a fixed-parity sector of the Hilbert space, but like the ordinary Clifford group only manages to be a 3-design.
Overall, the rest of this paper is structured as follows: In Section II we formally define Majorana strings, prove some basic properties, and discuss connections between operator weight, operator parity and the imposed superselection rules. We also highlight the relation to Pauli strings and the resulting Hilbert space representation. Next, in Section III we formally define the general (Majorana) Clifford group and identify the p-Cliffords as a subgroup, characterized in terms of orthogonal matrices over the field $\mathbb{F}_{2}$ and efficiently generated by braiding operators. In Section IV we discuss fermionic stabilizer states and how – with the help of an ancilla system – they can be embedded using p-Cliffords. Fourth, in Section V we discuss the problem of randomly sampling from the p-Clifford group, analyze its frame potential, and show that the p-Clifford group is a 3-design within sectors of fixed fermion parity. We conclude with an outlook and discussion of open problems in Section VI.
Example implementations of the algorithms explored in this work are available as a Julia Pluto.jl notebook. The notebook can be found in the supplemental materials, or can be provided on direct request.
@@ II Majorana Strings
To describe operator products of fundamental Majorana fermion modes and the superoperators that map between them, it is most convenient to use the language of vector spaces over the binary field $\mathbb{F}_{2}\equiv(\{0,1\},+,\cdot)$. This construction follows analogous (and equivalent) programs applied to ordinary Pauli strings [23].
@@ II.1 General Properties
Given $2n$ independent Majorana fermion operators $\chi_{i}$ satisfying (1), a general Hermitian Majorana string $\operatorname{\mu}(v)$ is defined (up to an overall sign) to be the normal-ordered product of said operators:
$\operatorname{\mu}(v)\coloneqq(i)^{v^{T}\operatorname{\operatorname{\omega}_{L}}v}\cdot\chi_{1}^{v_{1}}\chi_{2}^{v_{2}}\cdots\chi_{2n}^{v_{2n}},\quad v\in\mathbb{F}_{2}^{2n},$ (4)
@@
It is straightforward to show that this definition indeed satisfies $\operatorname{\mu}(v)^{\dagger}=\operatorname{\mu}(v)$ by reordering the operators after taking the adjoint. All expressions involving elements of the finite field $\mathbb{F}_{2}$ are understood to be evaluated mod 2. This also implies that complex phases with elements of $\mathbb{F}_{2}$ in the exponent don’t obey the usual algebraic rules [^1] The reason for this is the incompatibility between $\mathbb{F}_{2}$ having addition mod 2 and $(i)^{a}=(i)^{a+4}$ implying addition mod 4. The same issue also plagues the definition of qubit Pauli strings. and should therefore be treated with care.
With that we can define the equivalent of the Pauli group as the span of all (not necessarily Hermitian) Majorana strings:
@@
The group structure of $\operatorname{\mathcal{M}}_{2n}$ follows directly from the composition rule for two (Hermitian) Majorana strings:
@@
Here $\zeta(v,v^{\prime})$ is some non-trivial phase that arises partly due to the aforementioned algebraic quirks of the definition, and which can be shown to be of the form
@@
where we also identified the symmetric function
@@
Both expressions $\zeta(v,v^{\prime})$ and $f(v,v^{\prime})$ involve a new symmetric [^2] Analogous matrices over fields of higher order (corresponding to qudits with $d>2$ prime) are actually skew-symmetric, but for $\mathbb{F}_{2}$ this makes no difference. matrix $\operatorname{\omega}$ which arises in the derivation as the sum of $\operatorname{\operatorname{\omega}_{L}}$ and its transpose $\operatorname{\operatorname{\omega}_{U}}$:
@@
While (8) and (9) are nontrivial expressions, it is straightforward to argue (and explicitly prove) that
@@
since $\mu(v)$ is Hermitian with ±$\pm 1$ eigenvalues, and therefore satisfies $\mu(v)^{2}=I\equiv\mu(0)$. Furthermore, the overall phase simplifies dramatically (and becomes real) in the case of $v^{T}\operatorname{\omega}v^{\prime}=0$, which we will now show corresponds to $\operatorname{\mu}(v)$ commuting with $\operatorname{\mu}(v^{\prime})$.
@@
$\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})=(-1)^{v^{T}\operatorname{\omega}v^{\prime}}\cdot\operatorname{\mu}(v^{\prime})\operatorname{\mu}(v).$ (14)
The aforementioned expression $v^{T}\operatorname{\omega}v^{\prime}$ therefore indicates if two strings commute or anti-commute. And since any operator commutes with itself, it also has to satisfy
@@
defines a symplectic product, i.e. it is bilinear, alternating ($\left\langle v,v\right\rangle=0$) and non-degenerate [^3] By non-degenerate we mean that the matrix does not have zero eigenvalues. This should not be confused with the notion of degenerate eigenvalues in quantum mechanics.. $\operatorname{\omega}$ is therefore called a symplectic form.
@@
II.3 Operator Weight & Parity
As touched upon in Section I, the concept of parity is very important for systems composed of Majorana fermions. It does not only apply to the states living in the Hilbert space, but also to the operators acting on them. For Majorana strings we hence define the notions of (operator) weight and parity in the following way:
Definition 4 (Operator Weight & Parity).
@@ II.3 Operator Weight & Parity
  • The (operator) parity $\operatorname{p}(\operatorname{\mu}(v))\equiv\operatorname{p}(v)$ of a Majorana string $\operatorname{\mu}(v)$ can be defined as the string’s operator weight mod 2. Equivalently, the parity can be defined in terms of the dot product of the string’s vector representation with itself or $j$ (12) (over $\mathbb{F}_{2}$):
@@
As alluded to before, the observable distinguishing between operators and states of different parities is the aptly named parity operator
@@
which is uniquely defined up to an overall sign. Superselection then dictates that any Majorana string corresponding to a physical observable must commute with $(-1)^{F}$. However, using the general property
@@
one comes to the conclusion that $v^{T}\operatorname{\omega}j=v^{T}j=\operatorname{p}(v)$ for all $v\in\mathbb{F}_{2}^{2n}$ per (18), and thus the the commutation property (14) implies
@@
Hence only strings with even parity commute with $(-1)^{F}$ and can therefore be considered valid observables. This does however not imply that odd-parity strings are without use, as they can for example describe particle insertions or logical operators acting on stabilizer states. The latter is discussed in more detail in Section IV.3.
@@ II.4 Connection to Pauli Strings
$\operatorname{\pi}(v)=(i)^{v^{T}\operatorname{\operatorname{\eta}_{L}}v}\,X^{v_{1}}Z^{v_{2}}\otimes\cdots\otimes X^{v_{2n-1}}Z^{v_{2n}},$ (24)
where $X$ and $Z$ are the standard Pauli matrices and $\operatorname{\operatorname{\eta}_{L}}$ is the equivalent of $\operatorname{\operatorname{\omega}_{L}}$, here corresponding to the lower-triangular part of the matrix
@@
Analogous to $\operatorname{\omega}$, $\eta$ encodes the commutation structure of Pauli strings and hence also defines a symplectic product. We can therefore find an invertible linear transformation $W\in\operatorname{GL}(2n,\mathbb{F}_{2})$ that maps between these two symplectic forms:
@@
What $W$ therefore describes is the $\mathbb{F}_{2}$-representation of a Jordan-Wigner transformation between operator product strings of Majorana fermions, and tensor product strings of Pauli operators [24]:
@@
This is trivially true for $n=1$ since in that case $\operatorname{\omega}=\operatorname{\eta}$, which coincides with the natural identification $X\cong\chi_{1}$, $Z\cong\chi_{2}$. In general for our choice of basis, $W$ is of the form
@@
which is the upper-triangular part of the matrix complement of $\operatorname{\eta}$ (including the diagonal). To prove this, one first has to confirm that
$\operatorname{\eta}^{c}=(\operatorname{\eta}^{c})_{L}(\operatorname{\eta}^{c})_{U},$ (30)
which can be done e.g. by induction over $n$. Inserting $I=\operatorname{\eta}^{2}$ and applying (20) to the odd-parity rows/columns of $\operatorname{\eta}$ then yields
$\operatorname{\omega}=\operatorname{\eta}\operatorname{\eta}\operatorname{\omega}=\operatorname{\eta}\operatorname{\eta}^{c}=\operatorname{\eta}\,(\operatorname{\eta}^{c})_{L}(\operatorname{\eta}^{c})_{U}=(\operatorname{\eta}^{c})_{L}\operatorname{\eta}\,(\operatorname{\eta}^{c})_{U},$ (31)
where in the last step we also used the fact that $\operatorname{\eta}$ and $(\operatorname{\eta}^{c})_{L}$ commute. Additionally, (30) implies that one can also go the other way, as inserting $\operatorname{\omega}$ (which commutes with all matrices whose rows and columns have the same parity) into that expression results in
$\operatorname{\eta}=(\operatorname{\eta}^{c})_{L}\operatorname{\omega}\,(\operatorname{\eta}^{c})_{U}.$ (32)
In fact, one can prove (again using induction) that
@@
which means that $W$ both maps vector representations of Majorana strings to ones of Pauli strings, and vice versa. From the structure of (29) it also becomes apparent that low-weight strings in one basis are generally mapped to high-weight strings in the other, which illustrates the non-locality of the transformation. This is to be expected, as the most basic Jordan-Wigner transformations are used to assign local fermionic degrees of freedom to non-local topological domain walls, e.g. in the one-dimensional (bosonic) XY model [25].
The matrix $W$ does however not completely fix the corresponding Jordan-Wigner transformation, as it is agnostic with regard to specific sign choices for the operators. A canonical mapping with regards to our conventions is therefore provided in Appendix A. However, in the presence of a (bosonic) Hamiltonian $H$ it might be necessary to make a completely different choice for $W$ that complies with the structure of $H$, or even use a different type of transformation altogether [26]. The ramifications of this for our framework could be explored in future work.